40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 1 of 55  Not logged in ELOG logo
ID Date Author Type Category Subject
  2784   Fri Jul 30 16:55:38 2021 ranaElectronicsLaserDelay Line Freq Discriminators

I made a page in the ATF Wiki for Delay Line Frequency Discriminators. There is some prior work on these things, but these links are maybe a good starting point to see what the state-of-the-art is and whether our thing is better or not.

  2783   Tue Jul 20 19:15:43 2021 aaronDailyProgressLab Workfibers inspection

I started moving our fiber components to the rack-mounted box. I inspected and cleaned the tip of the fibers for the 50-50 fiber beamsplitter. Will work on some photos through the viewer of the analog microscope.

We need an uncoated APC to uncoated PC fiber patch cable to send a beam to our FC 1611. We have a coated patch cable, but should only use that for launching to free space.

I also searched for an appropriate replacement o-ring for the cantilever cryostat. I found the drawings and manual from IR labs, but they don't mention the size of the o-ring, and I didn't find something suitable in our supply. I'll order a new one, along with the patch cables and some panel mount DB9 adapters. 

  2782   Thu Jul 15 17:22:03 2021 aaronDailyProgressLab Worklab cleanup

[shruti, aaron]

We cleared out the desk drawers and the middle optics table today. This involved

  • sorting the papers in the draws into scratch paper / notes; manuals, datasheets, and MSDS; miscellaneous electronics diagrams and the like; papers and articles; black paper. We kept the manuals, datasheets, and MSDS, along with the papers and articles, and stored both in the small wooden cabinet along the S wall where we'd been keeping some books
  • Sorted the other materials (like cables, router, hard disks, other computer components, CDs, etc) into appropriately labeled clear plastic boxes, which we are storing under the workbench for now
  • Mounted the fiber breadboard box on the PSOMA rack, and put the associated connector panels inside the box
  • Decommissioned the cryo GeNS experiment
    • Gowned up and opened up the cryostat
    • Removed the Si wafer therein, and stored it in the desiccator cabinet
    • Detached the copper straps from the cold plate, and detached the wires coming from the electrical feedthrough at their kapton-wrapped pin connectors.
    • Wrapped the GeNS mount, including copper straps, periscope, substrate holder, ESD, schwarzschild, and the bottom of the cold plate in foil, and stored it in a clear plastic box. Also stored all in-vacuum screws, washers, and nuts in foil in the plastic box.
    • Removed the hot mirror from the cold shield and stored it in its original lens case in the optics cabinet
    • Closed up the cryostat and removed it from its mount. We stored the cryostat, the breadboard with an open center that it sits on, and its channel strut mounting hardware in the cage in the WB subbasement hallway.
    • Put the remaining channel strut materials making up the cryostat's mount with the rest of our channel strut hardware in cryo lab

We still need to clear the following heavier equipment from the desks. To move this equipment, we'll first need to sort through the miscellaneous objects on the workbench so we have somewhere to put everything. It will require at least two people to move the reference cavity safely.

  • reference cavity
  • gaston and spirou workstation computers, plus monitors and peripherals
  • stereo and speaker system

Other than that, we're ready for facilities to move out the desks and optics table. We should place an order for the new desks by tomorrow so they will arrive without too much awkward waiting.

  2781   Tue Jul 13 12:00:36 2021 aaronUpdateControl Systemventing cantilevers cryostat

[aaron, shruti]

morning

Shruti and I opened up the cantilevers cryostat. Vacuum wasn't fully vented despite the valve being fully open, due to the foil covering the open valve sealing against the flange while venting. The cryostat popped rather than lifted open. The damage is at least:

  • All three steel wires between the blade springs and optical platform snapped
  • both cantilevers broke and the mirrors detached from the cantilevers
  • there's a knick on the inner edge of the sealing surface we were opening, and a couple radial scratches visible.

Overall pretty disastrous. We'll have to investigate the full extent of the damage this afternoon, and first order of business will be cleaning out the broken Si fragments and evaluating the vacuum pressure.

Afternoon

We removed the silicon cantilever from the bottom of the cryostat with teflon-tipped foreceps, and stored them in a wafer casette. We also cleaned and regreased the o-ring with Krytox lubricant (we couldn't find the cryo lab's Apiezon N grease, must have been lent).

We pumped down the cryostat, but the roughing pump wasn't able to reach a low enough pressure to switch on the turbo. We also valved off the cryostat from the rest of the system, and pumped down on the hose + gauge + up-to-air valve (closed). The turbo was able to spin up to 90 krpm, but the pressure leveled out a several 100 uTorr. Pumpdown curves are in attachment 2.

The second pumpdown curve (pumping on just the hose and pressure gauge) suggests the pumping station needs some maintenance. I've seen this behavior before from a faulty KF flange connection, or when some condensation had built up in the roughing pump line. There is a procedure in the HiCube manual for flushing this line, I'll dig it up from my old elogs. However, the pumpdown on the cryostat suggests an even more severe leak in the cryostat itself, since the turbo wasn't even able to spin up and the curve appears leak-limited well above 1 torr. This is consistent with the visible damage to the mating surface in attachment 1. I suspect we need to send the cryostat to be reground and polished. crying

Attachment 1: EA078826-F058-4029-BE52-0A0534AF5180.jpeg
EA078826-F058-4029-BE52-0A0534AF5180.jpeg
Attachment 2: Screenshot_from_2021-07-13_16-25-28.png
Screenshot_from_2021-07-13_16-25-28.png
  2780   Mon Jul 12 15:34:30 2021 aaronLab InfrastructureDAQtemperature trend, cominaux apt

Attached is the most recent month of temperature trends for the lab, as measured by the particle counter. The lab temperature has been steady at 78 F for a couple weeks. I added a config file for this plot to our scripts repo under ndscope, so we can easily reproduce this plot.

While the particle counter has been logging, the AD590 temperature mon is again not logging. Possibly relatedly, when I try to apt-get install emacs on cominaux, it's complaining that the ligo repos 'couldn't be verified because the public key is not available.' I noticed that it was looking for the buster distribution in http://apt.ligo-wa.caltech.edu/debian, but the code currently lives in */debian/pool. Adding /pool/ (and commenting out the lines referring to stretch, which we are no longer using) let me upgrade cds-workstation and install emacs.

I did find some apparent bugs in the modbusIOC database file (CRYOXT.db), but none restored these ai channels. Haven't quite finished yet.

Update: I identified that only the slow ADC channel 15 was not logging by observing a 1 Hz sine from a function generator on the other channels. Unfortunately we use channel 15 for the AD590 temperature sensor. There was just an extra character in the epics record for that channel, removing it fixed the problem and the AD590 is again logging. I've updated with a new temperature trend that includes both sensors and the particle counts in attachment 2. I used dataviewer because ndscope doesn't support log axes, but also defined some new calc channels to record the log of particle counts... and upvoted the log-axis feature request for ndscope.

Attachment 1: Screenshot_from_2021-07-12_15-51-07.png
Screenshot_from_2021-07-12_15-51-07.png
Attachment 2: Screenshot_from_2021-07-13_10-52-50.png
Screenshot_from_2021-07-13_10-52-50.png
  2779   Mon Jul 12 14:28:38 2021 aaronElectronicsLab Workno noisy pickoffs for PDH signals

We've been using T-junctions to pick off PDH control and error signals, but sometimes we should have a high impedance, floating buffer between our current control loop and noisy or grounded devices like the ADC or oscilloscopes. I've added an SR560 between the PDH error signal and control signal pickoff points, and also used the 'aux out' connector on the back of the LB servo box as the control signal monitor point. I terminated aux out into 50 Ohm, with a T to the SR560 high impedance input.

Changes to the previous configuration are highlighted in attachment 1.

Attachment 1: 321F64CD-6EDA-4E4F-B9D5-0C403286E2AF.jpeg
321F64CD-6EDA-4E4F-B9D5-0C403286E2AF.jpeg
  2778   Thu Jul 8 19:10:59 2021 aaronDailyProgressLab WorkSetting changes

We should note that the LB box driving into 50 Ohm is current limited at its output (assuming we don't narrow the voltage window using the trim pots on the back panel). It can supply up to +- 20 mA, so if we see control voltages approaching 1 V we are nearing saturation of the servo controller.

Quote:

I also added a 50 ohm terminator in parallel to the 20 dB attenuator at the analog input modulation port of the ITC 502 current driver. This is to ensure a more or less accurate 20 dB attenuation of the control signal since the ITC 502 input has an impedance of 10 kOhm.

  2777   Thu Jul 8 14:51:08 2021 shrutiDailyProgressLab WorkSetting changes

Following our discussion earlier when we realized that the AC electronics of the 1811 may be saturating, since our RF power (at the mod freq of 33.59 MHz) is near or over 55 microW, I added an additional 10dB attenuator before the EOM. The total attenuation is now 40 dB. To compensate for this I removed the 10 dB attenuator at the input of the LB1005.

I also added a 50 ohm terminator in parallel to the 20 dB attenuator at the analog input modulation port of the ITC 502 current driver. This is to ensure a more or less accurate 20 dB attenuation of the control signal since the ITC 502 input has an impedance of 10 kOhm.

Everything seemed to lock once again when the gain on the LB1005 was increased from 5 to 6.

 

  2776   Thu Jul 8 09:50:41 2021 aaronDailyProgressNoise BudgetPSOMA noise budget, does it make sense?

I'm trying to make sense of these noise curves in relation to the free running noise I measured earlier with the three corner hat (3CH). The free running frequency noise of the S laser as measured by the three corner hat (attachment 6 from elog 2740) should be the same as the estimate off the free running noise using the PDH error signal and normalizing out the loop suppression ('open loop estimate' in attachment 2 of elog 2775). Here's what I'm noticing:

  • At low frequency (1 Hz - 100 Hz), the 3CH measurement shows the noise falling off like 1/f with a corner around 100 Hz. The PDH noise curve shows nearly flat noise or slightly increasing noise below 100 Hz, and is almost 10 MHz/rtHz compared to <10 kHz/rtHz on the 3CH. I'd expect the true free running laser noise to be closer to < 10 kHz/rtHz in this range.
  • At mid frequency, (100 Hz - 10 kHz), the PDH error signal noise curve is either itself noisy or contains many features, and falls off like 1/f. The 3CH curve is featured, but nearly flat at several 100 Hz/rtHz. The resolution bandwidth for the PDH error signal noise curve in this band seems to not allow sufficient averaging.
  • At high(ish) frequency (10 kHz - 100 kHz), both curves are nearly flat and with fewer features, but the noise measured by the PDH error signal is still 2 orders of magnitude larger than that measured by 3CH.

We should certainly try to resolve these discrepancies... perhaps we are seeing extra noise due to the electronics used for locking (the PDH measurement includes the LB box, analog RF electronics, a long DB9 cable to ADC that was observed to inject noise around 100 kHz only partially attenuated by our ferrite toroid, etc). There might also simply be a bug in the noise budget script, I'll check it out.

I agree that the more-than-1/f dependence in the transfer function from attachment 1 above seems fishy. It's above 10 kHz, so it can't be due to the influence of the temperature control loop.

Update:

I've moved Shruti's NoiseSpectra.ipynb script to the cryo_lab scripts repository, and pushed the version she uploaded earlier. I also added the data to cryo_lab/data/PDH/Noise with git lfs.

I found a couple bugs in the script, and made some modifications

  • We needed to compensate for a 10 dB attenuator between the PDH error point and the LB box error monitor, but instead compensated for 5 dB (temporary mixup of watts vs volts). Resulted in a UGF around 68 kHz (instead of 100 kHz), and phase margin of 31 degrees (instead of 16 degrees)
  • Defined some variables for the attenuation, frequencies, and open loop transfer function to avoid repeated references to the columns of the data file
  • Bug in the dBmtoV function. Was telling me 0 dBm is 1e-6 Vrms, but it should be 0.224 Vrms. P[\mathrm{W}] = 1[\mathrm{W}] * 10^{(x[\mathrm{dBm}] - 30) / 10} \implies \mathrm{V_{rms}}=\sqrt{50[\mathrm{Ohm}]*1[\mathrm{W}] * 10^{(x[\mathrm{dBm}] - 30) / 10}}
  • scipy.signal.freqs_zpk wants frequencies in rad/s, but the 'cavityPoleReverse' function was supplying them in Hz
    • Also, to get the appropriate DC behavior of this transfer function (that is, the gain at DC should be our V-to-Hz calibration), we must scale the 'k' of zpk by the angular frequency of the pole (or rather 1 / the zero's angular frequency, since we actually want the inverse of the cavity pole). Otherwise, the pole (zero) frequency enters the DC gain in
    • Relatedly, we are correcting for the cavity response by multiplying the noise spectrum in V/rtHz by a DC gain and a single real zero at the cavity pole... but shouldn't we instead model the cavity as a complex pole pair with some Q like the finesse, and apply the transfer function to power before going to V/rtHz? Maybe for a high finesse cavity these reduce to the same, but our cavity finesse is O(10) so this might matter near the cavity pole. I did not modify the script to use a pair of poles.

H(\omega) = k \frac{\prod_m(i\omega - z_m)}{\prod_n(i\omega-p_n)}\implies |k|= |G_\mathrm{DC}| \frac{\prod_n|p_n|}{\prod_m|z_m|}

 

Following these modifications, the noise at 10 kHz - 100 kHz is closer to that measured by 3CH. The updated open loop transfer function and noise curves are attached. I'm not chasing after the remaining discrepancy yet, since we expect that the PDH error signal was saturating during the attached measurement (see Shruti's elog from today for what we're doing about that).

Attachment 1: Noise.pdf
Noise.pdf
Attachment 2: OLTF_fit.pdf
OLTF_fit.pdf
  2775   Tue Jul 6 14:39:49 2021 shrutiDailyProgressNoise BudgetCalibrating PSOMA noise budget

[chris, aaron, shruti]

  • We (Chris, Shruti) noticed that the offset changes on ERC_MON_RATIO when the lights are turned off/ on despite having a long pass filter; it also changes on PDH_CTL_OUT when the slow loop is off. This is probably the main reason why we need to lock the slow controls to an offset of PDH_SET to get the brightest spot.

We (Aaron, Shruti)  re-measured the PDH error signal slope for calibration since the previous measurements were for the settings before the mode-matching was optimized.

drive parameters

A: pk-pk voltage (mV)  B: peak separation time (us) C: sideband crossing separation (us) D: difference in drive at sideband crossings (mV) E: cavity pole (MHz) = 33.59*2/C*A F: cavity response (mV/MHz) = A/E
1 kHz, 3 Vpp 500 13.6 99.2 480 4.6 109
5 kHz, 3 Vpp 478 2.96 26.1 480 3.8   126

The cavity pole has changed but the peak-peak voltage of the PDH error signal seems roughly the same as measured on Thursday before optimizing the mode-matching. It seems like the different temperature setting we are now at has changed the polarization of light entering the cavity; there is no half-wave plate in the path between the fiber launch and input coupler.

Initial crude noise calibration 

I used the above estimate of the cavity pole and response along with the data measured on Friday to obtain a calibrated noise spectrum (red curve in Attachment 2), then for data above 100 Hz I used the linear estimate of the open loop gain and roll-off shown in Attachment 1 to obtain the blue curve in Attachment 2.

All data and the jupyter notebooks are in Attachment 3

Attachment 1: OLTF_fit.pdf
OLTF_fit.pdf
Attachment 2: Noise.pdf
Noise.pdf
Attachment 3: initial_calibration_data.zip
  2774   Fri Jul 2 11:29:52 2021 shrutiDailyProgressNoise BudgetCalibrating PSOMA noise budget

Some observations

  • When I entered the lab the cavity was already locked with a very bright spot. The transmission was over 1000 counts, and I was previously seeing 600, which was after Aaron improved the mode-matching yesterday.
  • (Attachment 1)  It seems like the cavity was locked for almost 6 hrs with the temperature loop enabled before I entered.

Changes I made:

  • After tweaking around with the gain to improve the UGF of the loop, I set the gain to 5.5 from 4.7. Further details in next section.
  • I changed the gain on the PDA20CS transmission PD to be 40 dB, it was previously 70 dB. I also adjusted the waveplate in the curved optic transmission to send more light to the PD and less to the camera. Many camera pixels were saturated. So now the X1:OMA-ERC_MON_RATIO is around 4 when it was previously around 8, and the transmission X1:OMA-ERC_TRANS_MON_OUT16 is at 400 counts, previously at 1000.
  • I changed the KI value for the loop from 0.5 to 0.2 because I thought that there were very slow oscillations, but I don't think I noticed anything change when I did that, but I left it at 0.2.

Maximizing the UGF:

  • As I increased the gain on the LB1005, at around 5.5 I began to see a prominent broad peak at around 120 kHz in the noise spectra of the PDH error signal (Attachment 2). At roughly 5.7, I began seeing some saturation indicated by the presence of higher harmonics, which increased when I increased the gain further (Attachments 3 and 4).
  • (Attachment 3) This made it possible to achieve a max UGF of ~120 kHz with other settings remaining the same.
    • The offset from 0 dB to +3.16 dB   [8Jul21 edit: It should be 5 dB (voltage attenuated bt 3.16) since the attenuator was 10dB] is because of the presence of an attenuator at the input of the LB box, when measuring  PDH error signal/ LB error signal
    • Technically the units on the axis is dB and not dBm for the yellow trace, which is PDH error/ LB error similar to the measurements described earlier.
  • Not sure where the saturation (harmonics) is coming from.

 

Noise spectra

  • With the settings described here (from earlier), the uncalibrated noise spectrum is what is shown in Attachment 6. The data below and above 1 MHz were measured separately with different res BWs.
  • (Attachment 7) With the settings at present, I took several noise spectra measurements with different frequency ranges and res BWs and overlayed them.

 

I saved all the data from the Moku I measured today in the mokuliquidwb Google Drive.

Attachment 1: Screenshot_from_2021-07-02_11-32-06.png
Screenshot_from_2021-07-02_11-32-06.png
Attachment 2: MokuSpectrumAnalyzerData_20210702_141330_5p5_Screenshot.png
MokuSpectrumAnalyzerData_20210702_141330_5p5_Screenshot.png
Attachment 3: MokuSpectrumAnalyzerData_20210702_141402_5p7_Screenshot.png
MokuSpectrumAnalyzerData_20210702_141402_5p7_Screenshot.png
Attachment 4: MokuSpectrumAnalyzerData_20210702_141416_5p8_Screenshot.png
MokuSpectrumAnalyzerData_20210702_141416_5p8_Screenshot.png
Attachment 5: MokuFrequencyResponseAnalyzerData_20210702_141630_PG_Screenshot.png
MokuFrequencyResponseAnalyzerData_20210702_141630_PG_Screenshot.png
Attachment 6: Spectra20210622.pdf
Spectra20210622.pdf
Attachment 7: Spectra20210702.pdf
Spectra20210702.pdf
  2773   Thu Jul 1 13:39:12 2021 aaronDailyProgressNoise BudgetCalibrating PSOMA noise budget

[aaron, shruti]

I made a careful measurement of the cavity length with a high-precision rule, and got 582 mm. The FSR of our ring cavity is about 515 MHz.

I'm measuring the pk-pk voltage and frequency spacing of the PDH error signal to get the cavity response. I've converted the peak separation from seconds to Hz by taking the separation of the sideband zero crossings to be twice the modulation frequency (33.59 MHz), and assuming the triangle wave drive is a true linear ramp. The slope of the line connecting the peaks in the PDH error signal is a factor of 2 shallower than the derivative of the PDH error signal near zero (see this note from Tobin Fricke, and we also confirmed this numerically).

As a check, I've also recorded drive voltage at the sideband crossings. The control signal calibrated using the drive voltage should eventually give us the same noise estimate as the error signal calibrated with the cavity response.

drive parameters pk-pk voltage (mV) peak separation (time) sideband crossing separation peak separation (freq) difference in drive at sideband crossings (mV) cavity response (mV/MHz) cavity pole (MHz)
123 Hz, 0.3 Vpp 522 66 us 1.62 ms 2.74 MHz 124 381 1.37
1 kHz, 1 Vpp 496 3.68 us 72.6 us 3.41 MHz 140 291 1.70
5 kHz, 0.6 Vpp 486 1.76 us 33.6 us 3.52 MHz 198 276 1.76

We noticed some distortion of the error signal near either turning point of the current drive. We measured the power transmissivity of the input coupling mirror to be closer to 0.02 than 0.10 (by measuring the power before and after that mirror, with the cavity unlocked), probably because we have not set the polarization and the reflectivity is slightly higher than specified for the other polarization. With the lower transmissivity, the cavity pole is consistent with expectation.

We're inclined to believe the higher frequency measurements, since the low frequency one was made at low amplitude and thus had some distortion of the error signal close to where the drive turned around (and the 1/f frequency noise).

We did try to make the measurement at 10 kHz and 100 kHz, which should still be below the cavity pole... but didn't see a clean error signal.

Made a couple measurements of the noise spectra (PDH error signal, control signal) at various frequencies. Haven't grabbed them yet to upload.

We've noticed that the ratio of transmitted / reflected light (MON_RATIO) monotonically increases as we increase the PDH control setpoint for the temperature loop (that is, higher currents always increase this ratio). We noticed that the TRANS_MON and REFL_MON channels had some offset with the beam blocked (-300 and +15 counts respectively), but negating these offsets does not change the behavior.

Later, with the cavity locked I (aaron) adjusted the input steering mirrors to maximize MON_RATIO. The transmitted beam spot and PDH error signal both look cleaner. Moreover, after improving the mode matching MON_RATIO is no longer monotonic with PDH_SET. Perhaps that behavior was an artifact of some higher order mode resonances. Messed around with the combination of attenuators, gain, limits, etc. Before realignment and this work MON_RATIO was at most ~2.8; after it's reliably > 8. Should really check the DC responsivity of the photodiodes to determine if this is a reasonably good number, but the spot looks bright.

Also increased the limiter window on the LB box output to the full range... though that turned out to be only +- 5V, rather than +- 10V specified in the datasheet. Also, I'm driving the 10 kOhm input impedance current mod in directly, rather than T-ing off into a 50 Ohm terminator... not sure why this should be preferable? The output of the LB box is current limited into 50 Ohm (only supplies +- 20 mA, regardless of the voltage limiter), and with the attenuator I wouldn't expect reflections.

There's some 100 kHz oscillation present even with the loop off. It's somewhat affected by stressing the DB9 cable to the ADC, so perhaps adding some extra turns around our ferrite toroid would eliminate this.

  2772   Wed Jun 30 17:26:42 2021 aaronDailyProgressNoise BudgetCalibrating PSOMA noise budget

There is also a 'closed loop' way to make this measurement.

At the end of the day, we would like to know the coefficient from 'frequency to PDH error,' as well as that from 'control signal to frequency.' If we measure one and the open loop transfer function, we hae the other. We require either detailed knowledge of the cavity parameters, or an independent measure of the laser frequency. Fortunately, because we can drive the laser frequency, our independent frequency discriminator doesn't need to be as quiet as our cavity.

Earlier, I set up a 'delay line frequency discriminator' to measure the Teraxion laser frequency noise in a three corner hat with our two Rio lasers. The same discriminator can be used to measure the transfer function from the PDH error summing point to the laser frequency as measured by the delay line.

Alternatively, since we expect the PDH error signal response to frequency deviations to be flat, we could measure the beat note frequency with the moku phasemeter at a single frequency and do a 'lock in' measurement in post processing.

Attachment 1: 6AC149C5-9C0F-42E7-8DE6-AA1F5FB8D9AF.jpeg
6AC149C5-9C0F-42E7-8DE6-AA1F5FB8D9AF.jpeg
  2771   Wed Jun 30 14:38:49 2021 aaronDailyProgressNoise BudgetCalibrating PSOMA noise budget

I ended up debugging labutils/tektronix/tek_tools.py to grab this data from the scope.

I'm sweeping the laser current with a 0.5 Vpp triangle wave at 100 Hz attenuated by 10 dB. On the scope I'm measuring the PDH error signal and the triangle wave before the attenuator, triggering once only on the triangle wave. Data are in cryo_lab/data/PDH/210630_PDH_sweep.txt, and the analysis in cryo_lab/scripts/PDH_calibrate.ipynb does the following:

  1. trim the data to only look at one sweep through the PDH error signal
  2. Fits a triangle wave to the drive data, to clean up the x axis points
  3. Fits a nominal PDH error signal vs current modulation voltage to the observed PDH error signal vs nominal modulation voltage (given by the triangle wave fit)

This is not a great procedure. The PDH error signal has some features, and the data underconstrain the free parameters. The best fit is nowhere close to expectation (for example, estimates the cavity length at 19m, compared to about 0.5 m), and shouldn't be trusted. The variances for some of the fit parameters are large compared to the estimated values.

Attachment 1: 210630_PDH_fit.pdf
210630_PDH_fit.pdf
  2770   Tue Jun 29 14:13:53 2021 aaronDailyProgressNoise BudgetCalibrating PSOMA noise budget

We've been measuring the amplitude spectrum of our PDH error signal in V/rtHz. We'd like to calibrate that into Hz/rtHz of phase noise on our laser.

To measure the voltage-to-frequency responsivity of our laser driver and laser, we can sweep the current modulation input with a triangle wave while measuring the PDH error signal. Fitting the PDH error signal tells us the conversion from Volts to Hz (at wherever we inject the current modulation), the product of carrier and sideband powers,

To make the measurement, I'll

  1. Use a DS345 function generator to sweep with a 0.05 V amplitude triangle wave at 10 Hz at the current mod in port of our ITC 502 current driver
  2. While triggering on the triangle wave, record the PDH error signal on an oscilloscope
  3. Fit the PDH error signal to extract the cavity length (FSR), sideband and carrier power, responsivity of the laser frequency to excitations at mod in, cavity finesse, and some less physical voltage constants

Given the responsivity of laser frequency to excitations at mod in, we can calibrate the PDH control signal into units of Hz/rtHz. Alternatively, we could use the slope of the fitted PDH error signal and the FSR to calibrate the PDH error signal into units of Hz/rtHz.

I started making this measurement on the moku, but found some pernicious high frequency noise... possibly from the capacitance of the cable over to the electronics rack? We've seen it a few times before, and it produces a characteristic periodic structure in the FFT. Adding a 1.9 MHz lowpass filter after the function generator does not remove this noise, and it does not appear on the Tektronix oscilloscope monitoring the PDH error signal. Attachment 1 shows the measurement on moku.

Instead, I'll make the measurement by running a cable from the function generator to the TDS 3024B oscilloscope and record the traces over ethernet. The error signal still is a little more complicated than I'd like, probably due to the presence of higher order spatial modes... but I'll see what the fit says.

I ran an ethernet cable from the network switch on the PSOMA rack to the scope, and turned on GPIB by following instructions in the Tektronics manual. I tried dumping data from the scope over gpib using scripts I had been using at the 40m, now in the cryo_lab/scripts repo... but didn't have much luck. Chased around error messages for a while, and tried debugging tds3014.py, but didn't get it working.

python tds_dump.py 10.0.5.247 test.csv
Attachment 1: 8CD3FEA8-9EE1-4728-8639-A69206A8D098.png
8CD3FEA8-9EE1-4728-8639-A69206A8D098.png
  2769   Tue Jun 29 08:53:02 2021 ChrisHowToControl Systemacquiring slow channels

Here's how to add slow channels to the framebuilder DAQ:

  1. Edit /etc/advligorts/edc.ini to add the channels to the list
  2. Restart epics data collector: sudo systemctl restart rts-edc
  3. Restart framebuilder: sudo systemctl restart rts-daqd
  2768   Mon Jun 28 13:54:46 2021 aaronNotesLaserphase noise in diode lasers

Now that we are generating noise budgets, we'd like to compare our observed phase noise with that expected from the physical noise sources. Chris pointed us to some useful papers a while back, and I'm going to start documenting my further reading on the PSOMA wiki's noise budget page.

  2767   Mon Jun 28 13:16:07 2021 aaronDailyProgressControl Systemglitches

How did you get the framebuilder to recognize the slow channels defined in SoftIOC? I recall we ran into some problems with that last time.

I added your ndscope config file to the cryo_lab/scripts repo. Thanks!

  2766   Mon Jun 28 10:12:28 2021 shrutiDailyProgressControl Systemglitches, transfer function

The glitches seem to come from the slow controls output and temporarily result in an actuation voltage of +-10 V even when the limit for that channel was set to +-2 V in the PID script. This triggers the 'WIN' light and a beep from the ITC 502 while also losing the lock.

Aaron and I also had observed this last week.

 

[shruti, chris]

We were able to measure the loop transfer functions with the latest settings of KI=0.5 and PDH_CTL offset=2000cts on the slow control without being affected by the glitches too much. Chris also set up all the slow control channels so that they can be viewed with ndscope. In this we could also see that the transmission monitor sometimes glitched but the loop remained locked.

 

Attachment 2:

Updated measurement similar to elog 2759

Attachment 1: glitches.png
glitches.png
Attachment 2: OLTF.pdf
OLTF.pdf
Attachment 3: Setup.pdf
Setup.pdf
Attachment 4: Data.zip
  2765   Thu Jun 24 14:37:12 2021 aaronDailyProgressControl Systemslow temperature control, more transfer functions

I noticed the -B input of the LB servo was connected to the input of the Moku (via a 20 dB attenuator). I removed this connection. I also noticed that the PID control loop was still running in a tmux session ('temp_PID') on spirou; if you were trying to run the loop in a separate process, I could see that leading to some oscillation. It looked like the loop was only started once in the tmux session, which would have been yesterday before modifying the config file... so I suspect that's why the loop was oscillating. We should run this PID loop as a service on cominaux eventually.

For the last several days, we have been seen some ~MHz oscillations appear in the PDH error and control signals, even in the 'quiet' operation modes we found. Shruti noticed the signal disappeared from the control signal when we unplugged the cable connecting the control signal to the fast ADC. The DB9 cable is long, since it runs from the PSOMA rack to the cymac rack. We wrapped the DB9 cable 10 times around a ferrite torroid, and the oscillation disappeared!

We are also seeing some odd glitches in the temperature tune channel. The output of the Acromag DAC (as measured on moku oscilloscope, and indicated by the TEC hitting its window limit) has glitches where it hits the negative voltage limit (-10 V). The TEMP_TUNE epics channel is showing no such glitch, so it's almost certainly a hardware issue. We weren't able to reliably reproduce the glitch.

We played around with the servo settings to get a stable lock point. I didn't capture much data, since the glitches kept killing the lock.

  • LB box gain at 6.0, input offset at 5.08,
  • Temperature integrator gain set to K_I = 1, timestep at 0.1
  2764   Thu Jun 24 12:26:56 2021 shrutiDailyProgressControl Systemslow temperature control

I changed the error signal for the PID slow control to be the PDH control signal by modifying the file 'PIDConfig_SLDTemp.ini' in cryo_lab/scripts/temp_control on spirou.

The control signal does zero out when I actuate the loop, but the PID error signal shows strange oscillations when the control signal cable is plugged into the breakout board.

  2763   Wed Jun 23 17:53:21 2021 aaronDailyProgressControl Systemslow temperature control

[aaron, shruti, chris]

back to level 17

We swapped in a level 17 mixer, since the level 7 may have been saturating. We also switched from SLP-1.9+ to SLP-5+ after the mixer.

Slow temperature control

We continued setting up slow temperature control. Chris suggested nulling the LB servo offsets with the inputs terminated at 50 Ohms, instead of tuning them for the closed loop. With the new choice of offsets, we aren't seeing the 100 kHz oscillation.

In the locked state with temperature control handling the low frequency, we seem to not be locked exactly on resonance. The PDH error signal is the process for both current and temperature control. We used the ratio of cavity transmission to reflection (X1:OMA-ERC_TRANS_MON / X1:OMA-ERC_REFL_MON) to assess how close we are to resonance, and stepped around the PDH setpoint on the temperature control loop to maximize the ratio. However, at some point before the ratio is maximized the current loop loses lock. 

I suspect some stage of the PDH servo box is straying too far from its linear range. When we change the PDH setpoint on the temperature loop, it changes the DC level of the PDH error signal at the input of the LB servo box. Therefore, the signal entering the PI filter strays from zero. The deviation is about 20 mV at the LB box error monitor point when we lose lock (compared to +- 330 mV nominal range). I spent some time tweaking the input offset to adjust for the new setpoint, but wasn't very successful. I later realized we could feed the temperature PID setpoint to the -B input of the LB box, which would then adjust for the offset introduced by the temperature control loop. We'll try it tomorrow.

Finally, we updated the PID locking script to turn on only when enabled by X1:OMA-SLD_TEMP_EN, and when the cavity is locked (indicated by the trans mon channel X1:OMA-ERC_TRANS_MON_OUT16 reading > 50 counts). When the locking script is running, the logic is as follows

  • If not X1:OMA-SLD_TEMP_EN, then X1:OMA-SLD_TEMP_TUNE is set manually (the PID script does not change TEMP_TUNE)
  • If X1:OMA-SLD_TEMP_EN, then
    • If not X1:OMA-SLD_PDH_LOCK, then TEMP_TUNE is set to 0
    • If X1:OMA-SLD_PDH_LOCK, then the PID loop is engaged and TEMP_TUNE drives X1:OMA-SLD_PDH_SIG_OUT16 to X1:OMA-SLD_PDH_SET

Chris also added a GPIB controller to our ITC502, so eventually we can control TEMP_TUNE and read back TEMP_MON over GPIB instead of with Acromag channels. To use GPIB, local current control must be disabled, so this would require us to use the custom current driver or handle PDH locking at high frequency by phase modulating at the EOM. 

  2762   Wed Jun 23 11:36:49 2021 shrutiElectronicsLab Workswapping mixer

I just switched the mixer back to level-17 and changed the low pass filter to SLP-5+ (5 MHz corner)

Attachment 1: MixerandLPFupdate.pdf
MixerandLPFupdate.pdf
  2761   Tue Jun 22 14:35:40 2021 aaron, shrutiDailyProgressControl Systemslow temperature control

[aaron, shruti]

I observed an unexpected behavior this afternoon that I still can't explain. I managed to get the cavity locked using the LB box servo in LFGL mode. When I turned off the servo box by switching to 'lock off' mode, the cavity maintained lock and the PDH error signal was passed through to the current modulation point. Only the LB servo was driving the modulation point, and the temperature tuning was also off. My understanding from the LB manual is that in 'lock off' mode, no control signal is summed into the output signal... so why was the PDH error signal passed through?

Later, we started controlling the temperature of the laser diode using some slow epics channels.

  1. Lock the cavity using the LB box (in LFGL mode) to modulate the current
  2. Turn on the python PID loop. The script and configuration we used are in controls@spirou:~/cryo_lab/scripts/temp_control/.
    • Operated the temperature control loop at 5 Hz in pure integrator mode, with K_I = 0.1-0.5       
    • I have the PDH setpoint at 0, and checked that any DC offset on the ADC is small (< 10 counts)
  3. Next, we tuned the input and sweep offsets of the LB box to optimize our dynamic range. We need the temperature control loop on for this operation, to avoid railing the current modulation input at the laser driver
    • first, tuned the LB box input offset until the error monitor was centered on 0 V
    • Next, tuned the LB box sweep offset until the current control was centered on 0 V
  4. With the low frequency gain limited to 20 dB, the PDH error signal was wandering at a few Hz. We turned up the gain limit to 40 dB, though we could have increased the gain of the temperature PID until it took over at those frequencies.

We observed a 100 kHz oscillation in the noise spectrum after this procedure. We weren't able to change the oscillation by tuning the laser current (within 10 mA) or servo gain (while maintaining lock).

We measured the open loop transfer function of only the LB servo box (feeding back its output to -B), and didn't see a feature at 100 kHz or an oscillation in the noise spectrum. We measured the transfer function in both 'lock on' and 'LFGL' modes. We did observe a broad peak near 4.5 MHz in the noise measured at the LB error monitor (attachment 1, 2. The sharp peaks are artifacts from the Moku present even with no input connected).

Data are available in the ligo.wbridge google drive. Attachment 7 shows the broad 100 kHz oscillation on the PDH error signal in purple.

 

Attachment 3: Updated OLTFs with LB1005 measured separately

We measured the OLTF of the LB1005 servo by feeding back to itself with the setting 'LFGL' (Low Frequency Gain Limit) set to 40 dB, INT (pure integrator until it hits LFGL), and gain of 5.1 at 300 kHz.

 I (Shruti) think the gain setting was slightly different today which makes the green curve slightly higher in magnitude than the orange curve, but otherwise it seems to track it and do nothing strange. The orange curve is the TF of the LB1005 derived when the servo is used to lock the laser to the PSOMA cavity as in elog 2759. The phase for both measurements also seems to be the same up to 1 MHz.

Data for LB1005 in Attachment 4 and remaining data used for the plots in Attachments 3 and 4 of elog 2759.

 

Attachment 5: Measured OLTFs again with new settings

After these changes we measured the loop TFs again. It is strange that the phase of the open loop first increases and then decreases.

Also there is a weird dip at 80 kHz (right above the UGF) in both the Plant TF and the full open loop TF.

The LB1005 measured separately and in the full closed loop differ by ~4dB, the full loop setting resulting in the lower curve, even with the same gain setting. Otherwise the two run almost parallel, at least below 1 MHz.

The data for this is in Attachment 6.

Attachment 1: MokuSpectrumAnalyzerData_20210622_152905_1LB_2ctrl_LB1005only_Screenshot.png
MokuSpectrumAnalyzerData_20210622_152905_1LB_2ctrl_LB1005only_Screenshot.png
Attachment 2: MokuSpectrumAnalyzerData_20210622_153041_1LB_2ctrl_LB1005only_Screenshot.png
MokuSpectrumAnalyzerData_20210622_153041_1LB_2ctrl_LB1005only_Screenshot.png
Attachment 3: OLTF_wLBonly.pdf
OLTF_wLBonly.pdf
Attachment 4: LB1005TF.zip
Attachment 5: OLTF_wLB_20210622.pdf
OLTF_wLB_20210622.pdf
Attachment 6: 20210622LoopTFs.zip
Attachment 7: 65DBE9F2-84F8-4BA1-9A8B-4CF90D1ADE8A.jpeg
65DBE9F2-84F8-4BA1-9A8B-4CF90D1ADE8A.jpeg
  2760   Mon Jun 21 15:34:09 2021 aaronElectronicsLab Workswapping mixer

I swapped out the level 17 mixer in our PDH setup for a level 7 (ZFM-2-S+), and put it all in a box (attachment 1).

This let me remove the 10 dB attenuator before the LB box, though I had to add a 6 dB attenuator on the LO.

I measured the open loop transfer function again with the new mixer, and saw a 34 kHz UGF (attachment 2. I wasn't very careful to maximize the gain of the LB box). No correction was needed, since there is no longer an attenuator at either input of the LB box.

Attachment 1: A172449C-4B5D-47C1-B7FC-014B6CC28471.jpeg
A172449C-4B5D-47C1-B7FC-014B6CC28471.jpeg
Attachment 2: 796CBB56-3AA5-48D6-A344-F8239D9E605D.png
796CBB56-3AA5-48D6-A344-F8239D9E605D.png
  2759   Mon Jun 21 13:54:08 2021 shrutiDailyProgressPSOMAOpen Loop Transfer Functions

Having added the attenuator, A(s), at the input A of the LB1005 the loop algebra is changed slightly: Attachment 3 has the algebra and Attachment 4 helps with understanding the symbols. I have just considered this attenuator separately from the plant and servo.

Attachment 1: Open Loop TFs

  • The yellow curve is the actual open loop transfer function after subtracting 5dB in the magnitude of the ratio between the PDH error signal and the LB error signal to compensate for the 10dB attenuator at the input A of the LB box
  • The blue and orange magnitude curves were recorded directly from the Moku
  • The phase of the Math channel saved from the Moku seems to be a copy of the magnitude for all three OLTFs even though the screenshots seem to show a real phase (the data for this is saved in Attachment 3 and shown in the previous elog) so I re-calculated the phase but I'm not sure if it fully makes sense. (The calculation is in Attachment 4)

Attachment 2 is all the individual closed loop transfer functions that were recorded to calculate the open loop ones.

Attachment 3 has the data, settings, and screenshots recorded from the Moku to calculate OLTFs

Attachment 4 is the Jupyter notebook used to generate Attachments 1 and 2

Attachment 5 has the loop algebra and diagram

Attachment 6 is a diagram of the setup depicting the loop components

 

Quote:
 

\frac{V_\mathrm{PDH}}{B}=\frac{PG}{1-PG}\frac{A-B}{B}

\frac{V_\mathrm{LB,error}}{B}=\frac{1}{1-PG}\frac{A-B}{B}

\frac{V_\mathrm{control}}{B}=\frac{G}{1-PG}\frac{A-B}{B}

...

...

Indeed, we were able to eliminate the oscillations we had been seeing by adding a 10 dB attenuator between the PDH error signal and LB box input A, and changing the attenuator at the LB box output from 20 dB to 10 dB. [We also swapped out our ZFM-3H-S+ for ZFM-2H-S+, which has a 2 MHz low frequency cutoff compared to 50 kHz. Swapping mixers did not resolve the oscillation]

 

...

...

 

Attachment 1: OLTF.pdf
OLTF.pdf
Attachment 2: IndividualTFs.pdf
IndividualTFs.pdf
Attachment 3: OLTFs.zip
Attachment 4: LoopTFs.ipynb
{
 "cells": [
  {
   "cell_type": "code",
   "execution_count": null,
   "metadata": {},
   "outputs": [],
   "source": [
    "import numpy as np\n",
    "import matplotlib as mpl, matplotlib.pyplot as plt\n",
... 203 more lines ...
Attachment 5: NewLoopAlg.pdf
NewLoopAlg.pdf
Attachment 6: NewSetup.pdf
NewSetup.pdf
  2758   Fri Jun 18 10:42:08 2021 aaronDailyProgressPSOMAInvestigating the peaks

[shruti, aaron]

We messed around with the LB box today. Here's my attempt at a narrative from the wandering.

We are measuring transfer functions using the Moku and LB box error points.

  • Driving the 'B' channel on LB1005, and recording the transfer function to current control signal and LB error point (A-B). The ratio of these transfer functions gives us the open loop gain of the LB box PI filter, G.
  • Driving the 'B' channel on LB1005, and recording the transfer function to PDH error signal and current control signal. The ratio of these transfer functions gives us the open loop gain of the optoelectronic plant, P.
  • As a sanity check, also recording the transfer function from 'B' to both the LB error signal and PDH error signal. The ratio should give us PG.

\frac{V_\mathrm{PDH}}{B}=\frac{PG}{1-PG}\frac{A-B}{B}

\frac{V_\mathrm{LB,error}}{B}=\frac{1}{1-PG}\frac{A-B}{B}

\frac{V_\mathrm{control}}{B}=\frac{G}{1-PG}\frac{A-B}{B}

We first checked on an oscilloscope that we see no saturation at any of the monitor points (PDH error signal, LB error point, or current control point). We noted that the mean voltage at the LB box error point was reasonably close to zero (-2 mV). Then, ran a moku transfer function measurement with the following settings on the moku and LB box

  • Moku inputs; AC coupled inputs, 1 MOhm input impedance, 1Vpp range
  • Moku output: 50 mVpp drive for a 50 Ohm output load, with no offset
  • LB box: Input offset setting 5.0, PI corner at 0 Hz, 10 dB low frequency gain limit, proportional gain of 5.7.

We noticed that the control signal is much smaller than the PDH error signal, which seemed surprising at first blush since the LB box has a bunch of gain. However, the low frequency drifts of the control signal are much larger than those of the PDH or LB box error points (about 1 V drifts of the control compared to mV drifts of the error signal), which makes sense. And at high frequency the gain is falling off as expected.

I set the input voltage offset of the LB box by tuning the PDH error signal (on oscilloscope) to 0 V with the cavity locked; in retrospect, this isn't a rigorous procedure unless the temperature control is handling the DC error. With the input offset dial at 4.68 (-160 mV), the PDH error signal is centered at 0-2 mV a few mV drift over minute timescales. Before adjusting the input offset, the error signal was centered around -150 mV, which meant there actually was some clipping of the PDH signal! The input stage of the LB box saturates at +- 330 mV. The rms of the PDH error signal is 200 mV. Perhaps we should use a lower power mixer to avoid saturating the LB input (currently at level 17).

Indeed, we were able to eliminate the oscillations we had been seeing by adding a 10 dB attenuator between the PDH error signal and LB box input A, and changing the attenuator at the LB box output from 20 dB to 10 dB. [We also swapped out our ZFM-3H-S+ for ZFM-2H-S+, which has a 2 MHz low frequency cutoff compared to 50 kHz. Swapping mixers did not resolve the oscillation]

 THe LB box manual shows the electronic schematic for the servo in Figure 7. We again zero'ed out the inputs to the various amplifier stages by doing the following with both inputs (A, B) open and sweep range turned off:

  1. First, adjust the input offset such that the Aux Output reads 0V.
    • The dynamic range of the PI filter is +- 330 mV measured at V_error from Fig 7
    • I don't think the difference amplifier ever saturates before the filter amplifier, since the manual claims the input offset can null out DC voltages up to +- 10V.
  2. Then, adjust the sweep center such that the output also reads 0V
    • For the summation (last stage) amplifier, the max/min output voltages can be read by turning the sweep offset knob to the +- rails. The maximum output voltage is 1.43 V; the minimum is -1.72. We could trim the pots on the back of the box to make these appropriate to our current driver, and centered at 0 V

Attachment 2 shows the PDH signal and LB error point, so the math channel gives PG. Attachment 3 shows the PDH signal and control signal, so the math channel gives P. Attachment 4 shows the control signal and LB error point, so the math channel gives G.

Attachment 1: B09D5E30-08A4-45B1-934C-7D6E84FDBAAB.jpeg
B09D5E30-08A4-45B1-934C-7D6E84FDBAAB.jpeg
Attachment 2: 5368BBE2-F9C7-4DBC-B098-05363DB8FD20.png
5368BBE2-F9C7-4DBC-B098-05363DB8FD20.png
Attachment 3: 1A231013-D1A3-49BF-A5E7-0B4D9238E195.png
1A231013-D1A3-49BF-A5E7-0B4D9238E195.png
Attachment 4: 2B9714F6-0367-44F2-8E53-AF6E2C9ADFC0.png
2B9714F6-0367-44F2-8E53-AF6E2C9ADFC0.png
  2757   Thu Jun 17 12:14:33 2021 aaronDailyProgressLab WorkTemp tune not absolute

I measured the voltage from X1:OMA-S_TEMP_TUNE, and found caputing some value on that channel resulted in seemingly random voltages from that DAC channel. Modifying the epics record for X1:OMA-S_TEMP_TUNE to match the other analog output channels fixed the problem, namely commenting out the PREC, ASLO, and SCAN fields. The temp tune channel now reads the actual voltage from Acromag analog output channel 0.

I renamed the fast channels for clarity, and created some subsystems.

  • ERC = East Ring Cavity (anticipated a West cavity)
  • SLD = South Laser Diode (anticipating a North laser)
  • Also renamed the slow channels to match the new naming convention. X1:OMA-SLD_TEMP_[MON, TUNE]
Attachment 1: Screenshot_from_2021-06-17_12-54-07.png
Screenshot_from_2021-06-17_12-54-07.png
  2756   Thu Jun 17 10:35:42 2021 shrutiDailyProgress Temp tune not absolute

The 'TEMP TUNE' input is relative after all with a coefficient of 0.2 kOhms/Volt. Max input voltage \pm10 V. (Attachment 1)

The 'TEMP OUT' is absolute (as expected) with a coefficent 2 kOhms/Volt.

Attachment 1: Temp_Tune.pdf
Temp_Tune.pdf
  2755   Wed Jun 16 10:29:58 2021 shrutiDailyProgressPSOMAInvestigating the peaks

Attachments 1 and 2 are the same plots in elog 2751 but overlaid. They are the noise spectra measured at the output of the servo.

Attachment 1

Comparison between the LB1005 servo controller (yellow) and SR560 pre-amplifier(red) when the cavity is locked to the fundamental (TEM00) mode as seen on the monitor. Also shown in this plot (in blue) is the noise spectrum when the cavity is not locked to the fundamental, i.e., the temperature is detuned such that nothing was visible on the transmission camera monitor, but the SR560 servo is on.

Attachment 2

Noise spectra measured with the SR560 as the servo, but with different gains. All were measured while the cavity was locked to the TEM00 mode as seen on the transmission monitor.

 

 

Other updates:

I picked up 4 packages from Downs (three from ThorLabs and one from MiniCircuits)

Attachment 1: SameGainASD.pdf
SameGainASD.pdf
Attachment 2: DiffGainSR560.pdf
DiffGainSR560.pdf
  2754   Tue Jun 15 16:33:35 2021 aaronDailyProgressLab Workslow controls

[shruti, aaron]

Ran more cables to the ADC/DAC for PSOMA, and started setting up the soft channels to do slow control of the temperature setpoint.

  • Fast ADC channels 0-3 are mapped to the cavity PDH signal, control signal, reflection DC level, and transmission DC level respectively
  • Slow ADC channel 0 is mapped to the S laser driver temperature monitor
  • Slow DAC channel 0 is mapped to the S laser driver temperature tune

Beware, the temp tune setting seems to be an 'absolute' setpoint rather than a relative tuning on top of the setpoint from the front panel. We'll be testing this out in more detail.

  2753   Tue Jun 15 14:59:29 2021 shrutiLab InfrastructureDrawingsnew lab layouts

Discussed the lab layout with Aaron today and combined some ideas from both attachments in the previous elog

Attachment 1: lab_layout_plan.pdf
lab_layout_plan.pdf
  2752   Fri Jun 11 17:08:20 2021 ranaDailyProgressPSOMASpurious peaks

IT would be interesting to see an overlay of the ASD of the control spectra (SR560 low gain, SR560 high gain, LB1005 high and low gain). These are all directly comparable since they are at the actuator point. Might help us figure out where the peaks are coming from.

  2751   Fri Jun 11 08:58:44 2021 shrutiDailyProgressPSOMASpurious peaks

Yesterday Rana pointed out the peaks at 350 kHz and its harmonics suggesting that the loop was oscillating. He also added a RG850 long pass filter 3mm thick to the transmission PD. Despite the specs mentioning that the transmission at 1550nm was over 99%, this signal seems to have decreased (yellow trace in Attachment 5).

I also adjusted the servo gain at the SR 560 and found that the loop could 'lock' even with a gain of 5. I was not able to lock with gain of 2. Previously, the gain was set to 50. 

Investigating the spurious peaks

Attachments 1 and 2 are the PDH signal measured after the SR560 from its 600 Ohm output.

In Attachment 1, the cavity is PDH locked using servo (SR560) gain of 5. The peak at ~350 kHz and its harmonics are visible. In Attachment 2, I tuned the temperature away from the locking point while the servo is still active and the peaks are still visible. The same oscillations are also seen in Attachment 5. 

Attachment 3 and 4 are taken from the fast channels in the DAQ system. The gain at the SR560 was set to 50 for the reading in Attachment 3, and was set to 20 for the reading in Attachment 4. The gain setting of 50 seems to show better noise performance than 20 as seen on the digital system.

The purple trace and its persistence in Attachment 5 shows that the strength of this oscillating signal varies. This was measured with a gain setting of 5.

The oscillations were also seen when I locked the cavity with the LB1005 servo controller with gain=5 (lowest that allowed lock), PI-corner set to INT. Attachment 6 is the power spectrum of the output of the LB box when the cavity was locked.

Attachment 1: MokuSpectrumAnalyzerData_20210610_173153locked5_Screenshot.png
MokuSpectrumAnalyzerData_20210610_173153locked5_Screenshot.png
Attachment 2: MokuSpectrumAnalyzerData_20210610_173111unlocked5_Screenshot.png
MokuSpectrumAnalyzerData_20210610_173111unlocked5_Screenshot.png
Attachment 3: Locked_G50.pdf
Locked_G50.pdf
Attachment 4: Locked_G20.pdf
Locked_G20.pdf
Attachment 5: E6E3BD11-9678-466C-A2EE-FD848AB98E02.jpeg
E6E3BD11-9678-466C-A2EE-FD848AB98E02.jpeg
Attachment 6: MokuSpectrumAnalyzerData_20210611_121141_Screenshot.png
MokuSpectrumAnalyzerData_20210611_121141_Screenshot.png
  2750   Tue Jun 8 10:08:56 2021 aaronDailyProgressNoise BudgetPDH spectrum, transfer function

[aaron, shruti]

Using the channels we set up yesterday, we are measuring the PDH error signal and transfer functions.

  • first, just used dataviewer and diaggui to plot the channels. We're picking off the PDH signal before and after the SR560
  • Made a fresh MEDM screen for X1OMA, and populated it with one filter bank for the PDH signal before the SR560.
  • Used foton to apply a calibration gain from counts to volts.
    • Note that the foton picker is unhappy with python 3. We ran sitemap in the 'controls' conda env on spirou, which is running python 2.7, and were able to select the filter bank from medm screen.
    • Specifically, applied 10 V / 2^15 counts.
  • Attachment 1 is the power spectrum mentioned

Misc

  • Made an alias on cymac1 called 'rtreset' that stops all models, waits for them to stop, runs 'sudo rmmod x1iop', then starts all models. For some reason rtcds restart is not fully stopping x1iop, and this seems to avoid errors on restart.
Attachment 1: Screenshot_from_2021-06-08_13-05-16.png
Screenshot_from_2021-06-08_13-05-16.png
  2749   Mon Jun 7 13:23:33 2021 aaronDailyProgressLab Workmonitoring PDH signal on cymac

[aaron, shruti]

We would like to start recording our PDH error signal on cymac, both for noise budgeting and to provide slow feedback to the laser temperature.

Today we ran a couple cables from the PSOMA rack to the fast ADC/DAC boxes. Then, updated the x1oma model to record the N and S PDH error signals as _DQ channels. Next step is to define some softIOC channels for the PID controller. Attachment 1 shows the status lights after modifying the x1oma model, and disabling the models for older Cryo Lab experiments.

Attachment 1: status.png
status.png
  2748   Thu Jun 3 15:02:13 2021 aaronComputingstuff happensrestart rts-edc

We got a new HEPA FFU yesterday. I wanted to see if the particle count has dropped as a result, but the minute and longer trends are not available... apparently no trends have been logged for about three weeks. I see from systemctl that the rts-edc (realtime system EPICS data concentrator) loaded but failed to run. Restarting that service eventually restored the frames. We should get an email when this happens.

 

  2747   Wed Jun 2 15:42:20 2021 aaronLab InfrastructureDrawingsnew lab layouts

We are about to purchase a new clean enclosure for the PSOMA table, and while we've got facilities modifying the sprinklers would like to consider alternative floor plans for the lab. Here are two possible directions to go.

Attachment 1: We've removed the cryo Qs optics table, and moved the workbench to the center of the room. Desks are now located at the N and S walls of the room, and PSOMA table is close to the same location. Other cabinets are mostly unchanged, but I bumped the cymac rack over to improve our access to those cables.

Attachment 2: A more radical proposal. Move the workbench to the W wall to put a work area near the sink and decongest that corner. Move the vacuum cabinet a little north and optics cabinet a little south, making the optics a little easier to access. The wire rack moves to the S wall, since it doesn't have swinging doors and could overlap the lab's air intake slightly without blocking anything. PSOMA table is centered in the room, letting us consolidate the electronics racks and associated cables to one part of the room. I like this layout, but it might be tricky to move so many cabinets. 

 

For reference, the existing floor plan is approximately the same as Johannes' diagram from 2016

Update: Labelled the PSOMA optical table with dimensions. Other items in the lab are approximately to scale. I quite prefer the second option with the PSOMA table in the center of the room, as in attachment 2. 

Attachment 1: Canvas_1.pdf
Canvas_1.pdf
Attachment 2: Canvas_2.pdf
Canvas_2.pdf
  2746   Fri May 21 14:15:21 2021 aaronDailyProgressLasercurrent noise of custom current drivers

I'm measuring the noise of the cryo cavs current drivers. They are labelled with D1500207, and their traveler numbers are S1600247 for the W driver and S1600248 for the E driver.

First, I used a BNC breakout and multimeter to measure the DC voltage across a 20 Ohm resistor, while the interlock pins were connected to 50 Ohms. I found

  • both drivers have the same scaling from 'dial setting to current,' and can provide up to ~175 mA to the diode
    • The coarse knob has a dial with 50 ticks per turn, and can be turned up to 10 times. Each tick changes the current by 400 uA, so changing the number above the dial by 1 (ie, turning the knob once by 360 degrees) changes the current by 20 mA.
    • The fine knob also has 50 ticks per turn for 10 turns, but is 10 times more sensitive (40 uA / tick; 2 mA / turn).
    • There is about 0.5 mA supplied even with both current knobs set to '0'
    • There is some current limiter than prevents the current from changing when the coarse knob is turned from 9 to 10; in fact the current is no longer monotonic above setting 9 on the coarse knob
  • The low frequency current monitor of both drivers reports 20 mA / V.
    • If I send the LF current mon BNC to a non-floating oscilloscope (or moku), the driver cannot maintain its negative supply voltage and trips off. I was a bit surprised by this, since I didn't observe the current draw at the power supply changing -- maybe I missed it, or the ground reference near the output monitor BNC has some resistance between it and ground from the supply.
  • Disconnecting the interlock does not turn off the current. Effectively, there is no interlock for the laser when using these drivers.

 

Next, I wanted to measure the resistance of the diode at some usual temperature and current setpoints. I was overloading my SR560 buffer at DC, so instead I'm using a lock in amplifier to make the measurement at AC.

  1. First, I'm checking the responsivity of the current to drives at the HF modulation input BNC. For this, I'm using the SR554 since it has a very high common mode rejection ratio. Every other simple configuration I tried results in a ground loop that overloads the driver, causing it to trip off.
    • Attachment 1 is the measurement setup in this manner, and attachment 2 is the result on the lock in amplifier. The implied responsivity is ~1.1 mA/V
  2. I was stumped for a while on how to measure the resistance of the diode itself, rather than the current noise across it. What device has high input impedance, high common mode rejection, and a floating input relative to its output?
    • In the end, I think the Fluke multimeter in 'AC voltmeter' mode may be the best answer. Still, I was hoping for something that could help me plot diode resistance vs bias current / temperature so I looked into the other options below. Why is this simple enough that the multimeter can do it, yet seemingly so obscure that no SRS instrument can?
      • It would be great to have a Busby low noise box with an instrument amplifier instead of LT1128, so the input would be truly floating relative to the output!
    • Directly measuring the voltage across a dummy load ('laser diode') with an oscilloscope leads to short, since the shield of the BNC on the scope is tied to ground but the driver floats the 'laser diode' around 12 VDC. The driver overloads and turns off.
    • The SR560 preamplifier, SR830 lock in amplifier, and SR785 spectrum analyzer all have floating inputs. With AC coupling the DC voltage across the diode won't overload their amplifiers. However, none has enough common mode rejection to handle the 12 VDC floating 'diode,' so the inputs overload anyway. All would require some additional preamplifier to reject the common mode.
    • The SR554 transformer amplifier has excellent common mode rejection, but low (0.5 Ohm) input impedance. SR554 is fine for measuring the current noise of the drivers, but can't be used to measure the resistance of the laser diode. The transformer shunts almost any load parallel to its input -- instead of measuring the voltage across the load, it is directly measuring the supplied current and 'transforming' it to a voltage.
    • The busby box has high input impedance and common mode rejection, but it achieves this by passing the common mode to its output (the low side of its 'floating' input is the same as the low side of its floating output, and both are tied to the box chassis). When I use the Busby to drive an oscilloscope, there is a ground loop and the laser driver trips off. On the other hand, since it is a voltage source with 20 Ohm output impedance, it's not ideal for driving the SR554... but I think it could for small signals.
      • The LT1128 is operated with G=100 using a R=20 Ohm pulldown resistor and R=2 kOhm feedback resistor. Since there's a 20 Ohm buffer at the output, a ~V output would become O(50 mA), well above the 10 mA supply current limit of the op amp. But for a small signal O(30 mV) out of the AC-coupled Busby, the op amp would only have to supply O(1 mA)... maybe that could work. Since the LT1128 has protection against short-circuiting the output, I gave it a try and it seemed to work... but is very sensitive to capacitive coupling, like changing cable lengths, twists, or my touching the Busby box.
      • Driving the HF input with a sine of known amplitude at 200 Hz and applying the LD current across a known dummy load, I measured the LD current responsivity to HF excitation voltage at A_\mathrm{HF}=0.1 \mathrm{mA/V} (see attachment 3).

\frac{V_\mathrm{response}}{V_\mathrm{excitation}}=\frac{V_\mathrm{response}}{V_\mathrm{excitation}}\frac{1[V_\mathrm{554}]}{500 [V_\mathrm{response}]}\frac{20 [V_\mathrm{Busby}]}{0.5 [V_\mathrm{554}]}\frac{1[V_\mathrm{LD}]}{100 [V_\mathrm{Busby}]}\frac{R_\mathrm{sense}[I_\mathrm{LD}]}{1[V_\mathrm{LD}]}\frac{1[V_\mathrm{excitation}]}{A_\mathrm{HF}[I_\mathrm{LD}]} = 8\times10^{-4}R_\mathrm{sense}/A_\mathrm{HF}

If the Busby-to-SR554 system doesn't give the same responsivity as SR554 or Busby-to-flukemeter, I can't trust it to tell me the correct resistance of the laser diode. Might try the measurement again to see if I've missed something (it's not the lockin input coupling or filter settings visible in the photos, I tried), but if there's another approach I'm interested in hearing.

Attachment 1: 5012D147-8B13-478C-BB83-57A71A2B304A.jpeg
5012D147-8B13-478C-BB83-57A71A2B304A.jpeg
Attachment 2: 467DEC59-01B1-40CD-9F4E-4D9D87183A78.jpeg
467DEC59-01B1-40CD-9F4E-4D9D87183A78.jpeg
Attachment 3: ABC2A51B-6726-4F9C-AD88-551ED7C90EEF.jpeg
ABC2A51B-6726-4F9C-AD88-551ED7C90EEF.jpeg
  2745   Thu May 20 13:32:11 2021 aaronComputingstuff happenswhere's the x1cry medm master?

I went digging for what TEC settings Johannes had been using for those lasers, and what x1cry channels I might use to measure their PV curve to get a current noise.

Strangely, I can't find the x1cry medm master files. I recently copied over these files into the x1oma medm directory, then made some modifications (rm'ed some files unnecessary for x1oma). I'm seeing those modifications reflected in the medm/x1cry/master directory, as well as medm/x1oma/master. Here's my bash history

cp -r x1cry/master x1oma/master
cd x1oma/master
rm crymaster_old.adl
rm crymaster_orig.adl
rm crymaster_BAK.adl
mv crymaster.adl omamaster.adl
rm crymaster_setup_BAK.adl

I might not have actually lost data, since it looks like x1cry/master contained the actual master crymaster.adl, plus several backups. However, any changes I make in x1cry/master (eg, renaming the file) are reflected not only in x1oma/master, but also in the backups in medm/archives/x1cry_*/master. The files I'm working on are not linked, and the archives certainly shouldn't be affected by updates to the main files, so I don't know why this should happen.

Ah, I didn't realize x1cry/master was linked to some other directory. Now x1oma/master is linked to its own directory.

  2744   Thu May 20 12:50:56 2021 aaronDailyProgressLaserdelay line frequency discriminator

Now measuring the Teraxion laser beating with the Rio lasers for cryo cavs experiment. The E Rio laser is SN 5597; the W Rio laser is SN 5601.

I'm testing the WX beat note first, using the configuration attached to the above post. The Teraxion laser is sending 262 uW to our 1611.

laser pair beat note power TEC setpoint lower null LD current lower null frequency lower null DC voltage upper null LD current upper null frequency upper null DC voltage LD current at half fringe frequency at half fringe RMS voltage at half fringe frequency at end of measurement measurement timestamp
WX 442 uW 9.691 kOhm 107 mA 75.0 MHz -16.0 V 67.2 mA 203.3 MHz 0.9 V  88.0 mA 158.5 MHz 0.1 V   Thu May 20 17:00:33 2021
EX 449 uW 14.257 kOhm 89.4 mA 75.3 MHz -12.8 V 144.2 mA 204.0 MHz 2.8 V 124.6 mA 159.0 MHz 0.1 V 158.5 Thu May 20 18:01:48 2021
EW                          

Notes:

  • I'm recording the LD currents as those reported by the 'coarse' and 'fine' knobs of the custom drivers, assuming the actual LD current is 2x the reading from the knobs.
  • lower and upper null DC volltages are reported as 100x the voltage measured by the moku spectrum analyzer picked off from just before the AC coupled SR560
  • Instead of DC coupled SR560 with less gain, I've set the SR560 to AC coupled mode with G=100. Same LP at 30 kHz
  • I'm seeing some sidebands on beat notes involving the E laser. must not have a good operating point.

 

Current noise

 

  2743   Fri May 14 18:43:25 2021 aaronDailyProgressLaserdelay line frequency discriminator

I set up the attached configuration. The cryo cavs lasers are driven by the custom current drivers. I only have one remaining TEC controller (the other cryo cavs controller was moved to PSOMA rack). To drive the other TEC, I've moved the ITC510 from PSOMA rack to cryo cavs rack, and am using it to control the East TEC. Our TED 200 C is controlling the West laser TEC.

Attachment 1: 994ECC7F-C5FE-4E7B-8F19-B1C118DCE6CE.png
994ECC7F-C5FE-4E7B-8F19-B1C118DCE6CE.png
  2742   Fri May 14 13:27:42 2021 aaronLab InfrastructureGeneralwhat's up with nodus proxy errors?

 I'm getting frequent proxy errors when uploading pdfs to the elog. Started yesterday, continuing today. Error message below.

Proxy Error

The proxy server received an invalid response from an upstream server.
The proxy server could not handle the request GET /Cryo_Lab/.

Reason: Error reading from remote server

  2741   Thu May 13 19:09:16 2021 ranaDailyProgressLaserdelay line frequency discriminator

beyond ~70 Mhz, there will be some frequency dependence due to the frequency dependent loss in the cables. Perhaps also in the mixers and amplifiers. 

I suggest we just stick to the 10-20% cal uncertainty for now. If this setup seems good, however, we can rebuild it in a more robust fashio by ordering parts that are smooth out to 500 MHz.

  2740   Thu May 13 16:31:40 2021 aaronDailyProgressLaserdelay line frequency discriminator

I'm repeating the above measurement with a shorter cable (122 cm). The path length difference is now ~107 cm, which means the distance between adjacent nulls should be 93 MHz. If the beat note wanders by less than about 16 MHz, we should be linear enough. I decided against a shorter cable to ensure I'm within the 250 MHz bandwidth of the moku spectrum analyzer during calibration.

Calibrating the new delay line box with the marconi:

  • The first null is at 74.93 MHz. Mean DC output is -1.15 V.
  • The second null is at 110.6 MHz... odd, the FSR seems almost unchanged? Am I missing something?
  • There is a third null at 203.9 MHz (-.05 VDC). The spacing between 113.8 MHz and 203.3 MHz is closer to my expectation, and I was observing some odd steps in the signal near 75 MHz, so I'm going to use 114 and 203 MHz as my two nulls.
  • That puts the half fringe at 158.6 MHz. At the half fringe, the DC voltage is 0.80 V, and Vpp is 4.5 mV for a 100 kHz deviation FM at 1 kHz.
    • Although, I found that at 150.8 MHz the Vpp is 5.1 mV... call it a 10% calibration uncertainty

This time, I am recording the output from the delay line discriminator on moku while measuring the spectra with SR785. Should provide a check at the end that everything makes sense.

Laser pair beat note power TEC setpoint lower null frequency lower null DC voltage lower null LD current upper null frequency upper null DC voltage upper null LD current frequency at half fringe voltage at half fringe RMS voltage at half fringe Hz/V calibration Frequency at end  timestamp
NS 414 uW N: 9.692 kOhm
; S: 9.515 kOhm
110.7 MHz -1.90
 V
N: 104.7; S: 119.85 mA 203.7 MHz 0.03 V 120.61 mA 156.5 MHz -0.99 V 3 mV     Thu May 13 17:50:18 2021
NX 598 uW N: 9.692 kOhm 110.1
 MHz
-2.81 V 113.1 mA 203.3 MHz 0.26 V 114.2 154.6 MHz -1.50
 V
3 mV   144.8 MHz Thu May 13 18:40:15 2021
SX 369 uW S: 9.515 kOhm 111.3 MHz -1.98 V 121.38 mA 203.3 mHz 0.04 V 122.17 mA 156.6 MHz -1.04 V 3 mV     Thu May 13 19:21:12 2021

 

Due to the drift and small current change, the above aren't a great measurement of 'Hz per Amp' for the laser diodes. For example, if I assume the change in beat note frequency between the upper and lower null measurement of SX beat is due to changing the laser current of the S laser, then use that to estimate the expected change in beat note frequency of the NS beat going from the lower to upper null LD current, I'm off by 47 MHz compared to the observed change in beat note frequency between NS lower and upper null.

Here's a new calibration for both lasers, based on the NX and SX beat notes (up to sign):

laser frequency 1 LD current 1 frequency 2 LD current 2 f vs I
N 237.3 MHx 115.4 mA -237.7 MHz 103.9 mA 41 MHz/mA
S 238.7 MHz 122.49 mA -230.5 MHz 118.68 mA 123 MHz/mA

This still seems implausible to me. Our HP 8560E is a wider band spectrum analyzer, but it's doing some weird mode hop thing. I'll use the above for now. At least I have observed the S laser to have a larger frequency deviation per mA across a few measurements, and it's also consistent with my observing larger frequency drifts when the S laser is involved in the beat note.

The current noise measurement of the ITC510 is still too high..... attached is the three corner hat measurement from the SR785. I'll redo the current noise again and get the plot readable tomorrow.

Update: Oh, I'm just missing the G=100 from SR560 on the current noise measurement. With that factor, the ITC510 I below the beat note spectra. Will update the plots today.

Update: plots attached, now removing the filter from SR560. All attached plots are from the measurements by SR785.

Attachment 1: DL_NX_SPSR.pdf
DL_NX_SPSR.pdf
Attachment 2: DL_SX_SPSR.pdf
DL_SX_SPSR.pdf
Attachment 3: DL_NS_SPSR.pdf
DL_NS_SPSR.pdf
Attachment 4: DL_3CH_X_SPSR.pdf
DL_3CH_X_SPSR.pdf
Attachment 5: DL_3CH_N_SPSR.pdf
DL_3CH_N_SPSR.pdf
Attachment 6: DL_3CH_S_SPSR.pdf
DL_3CH_S_SPSR.pdf
  2739   Wed May 12 16:52:08 2021 aaronDailyProgressLaserdelay line frequency discriminator

I'm repeating the measurement above, this time with the delay line interferometer in a box. The cable length is slightly different -- I was using a 3 m BNC cable for the delay line before, and today made my own ~3 m SMA cable. The RF power level is also somewhat reduced, since I moved the attenuator before the splitter to avoid having to open the box and attenuate one leg of the interferometer. SR560 settings are same as yesterday.

I'm also calibrating with the marconi, rather than stepping around the laser current to null the beat note signal. The interferometer doesn't change between measurements.

  • At 108.9 MHz with a 10 dBm carrier, the mean DC output is -1.38 V
    • notice some distortion at the 'null' frequency, and can't completely null the signal
  • At 71.7 MHz with a 10 dBm carrier, the mean DC output is -0.04 V
  • That puts the half fringe at 90.3 MHz. With a 10 dBm carrier at the half fringe and modulation off, the rms voltage is 500 uV. With a 400 kHz deviation, 1 kHz FM on the carrier, the rms voltage is 12.1 mV (33.2 pk-pk). At 800 kHz deviation, the output of the discriminator is 24 mVrms, 66.6 mVpp. At
laser pair beat note power Rio laser TEC setpoint lower null frequency lower null DC voltage lower null LD current upper null frequency upper null DC voltage upper null LD current voltage at half fringe frequency at half fringe RMS voltage at half fringe Hz/V calibration frequency at end of measurement Timestamp during measurement
Rio N x Teraxion 452 uW 9.548 kOhm 71.7 MHz .03 V 96.1 mA 108.9 MHz -2.24 V 96.8 mA -1.04 V 90.9 MHz 0.02 V   99.0 MHz Wed May 12 17:56:28 2021; attachment 2
Rio S x Teraxion 365 uW 9.515 kOhm 70.1 MHz -.01 V 119.65 mA 108.3 MHz -1.67 V 119.33 mA -1.0 V 90.5 MHz -0.9 V   76.6 MHz Wed May 12 18:53:01 2021, attachment 3
Rio N x S 405 uW N: 9.691 kOhm; S: 9.515 kOhm 71.0 .01 V N: 104.7 mA; S: 119.26 mA 109.4 -1.88 V N: 104.6; S: 119.54 mA -1.0 V 90.6
 MHz
15 mV   95.9 Wed May 12 19:39:36 2021, attachment 4

Obviously, lots of drift throughout these measurements. I'll analyze but not much trust these results. One workaround to deal with the drift would be to increase the FSR by shortening the delay line, at the expense of some signal. Could also first lock the laser to the cavity, then make the same measurement (or even with a longer delay line).

I also noticed the effect of asymmetry splitter-delayline-mixer IFO by some DC offset at the bright/dark fringes.

Attachment 1: 3878D102-C5BB-46AE-B90B-540CFDE3100B.jpeg
3878D102-C5BB-46AE-B90B-540CFDE3100B.jpeg
Attachment 2: SPSR785_NX_12-05-2021_173959.pdf
SPSR785_NX_12-05-2021_173959.pdf
Attachment 3: SPSR785_SX_12-05-2021_183434.pdf
SPSR785_SX_12-05-2021_183434.pdf
Attachment 4: SPSR785_NS_12-05-2021_191513.pdf
SPSR785_NS_12-05-2021_191513.pdf
  2738   Wed May 12 15:35:09 2021 aaronDailyProgressLaserdelay line in a box

I put the splitter, mixer, and delay line in a box. The box is a little smaller than would have been ideal so I'm not completely satisfied with the strain relief, and one of the SMAs rotates due to the locking washer being overtightened... but I think it'll do. Innards attached,

  • parts are
    • mixer: ZFM-3-S+
    • splitter: ZFRSC-42-S+
    • lowpass: SLP-1.9+
Attachment 1: 69A8E6A2-7F85-464E-BDAA-8C6BBD2910B5.jpeg
69A8E6A2-7F85-464E-BDAA-8C6BBD2910B5.jpeg
Attachment 2: 1D05A31E-7DBC-448B-A2F4-A8805EFAC393.jpeg
1D05A31E-7DBC-448B-A2F4-A8805EFAC393.jpeg
Attachment 3: 25C2DD1E-7776-491E-BBCE-5F2D1EA935E3.jpeg
25C2DD1E-7776-491E-BBCE-5F2D1EA935E3.jpeg
  2737   Tue May 11 14:37:48 2021 aaronDailyProgressLaserdelay line frequency discriminator, current noise measurement

I'm repeating this measurement, this time recording the current noise of the commercial driver before each measurement for reference. Will update tomorrow with additional measurements, plots.

Delay line frequency discriminator procedure

  1. Set up the delay line frequency discriminator as before (see attachment 1)
  2. Find a low noise operating point (see below) for the Rio laser(s)
  3. Calibrate the delay line by tuning to the dark and bright fringes around 80 MHz and recording the DC voltage of the LF (after whatever amplification is applied)
  4. Measure the current noise of the commercial driver as below [could also be done completely independent of these steps]
  5. Tune the beat note frequency to the half fringe, and record the beat note spectrum.
  6. Repeat for each pair of beat notes.
  7. When done, turn off RF amplifier (first disconnect input, then turn off DC power) and SR560, and remember to plug in the SR560 power line.

The measurement parameters for delay line frequency discriminator are:

  • Pickoff RF from 1611 to Moku (50 Ohm AC coupled spectrum analyzer) using
  • RF amplifier ZHL-1A (powered by +24 V from GPS3030D) to splitter
  • Level 7 mixer ZFM-3-S+ (0.04-400 MHz)
  • LF from mixer through SMA-T with one end terminated at 50 Ohms, the other to SLP-1.9+ then SR560 channel A
  • SR560 settings: battery powered, DC coupled, 30 kHz LP (6 dB/oct), low noise mode, G=10
laser pair Beat note power Rio N laser TEC setpoint lower null frequency lower null DC voltage lower null LD current upper null frequency upper null DC voltage Upper null LD current voltage at half fringe frequency at half fringe RMS voltage at half fringe Hz/V calibration start time of measurement drift during measurement notes
Rio N x Teraxion 440 uV 9.548 kOhm 63.5 MHz .18 V 95.5 mA 101.1 MHz -3.28 V 96.2 mA -.95 V 82 MHz 35 mV   Tue May 11 17:26:42 2021   1dB attenuation on LO

 

Finding a 'low noise' operating point

According to the datasheet (PSOMA wiki 'documents' page), the lasers come with a recommended temperature setpoint (T_set on the datasheet). This setpoint may either lie on the upper hysteresis branch, or lie outside of the hysteresis region... but according to our datasheet, our lasers' T_set lie in the upper hysteresis branch. If we observe spectral distortions, this indicates the temperature is outside the recommended range or operating on the lower arm of the hysteresis. To achieve a low noise operating point

  1. Ensure the laser is properly mounted on the PCB with drivers and control loops connected (but off)
  2. activate the thermal control loop, and let the internal temperature stabilize
  3. Apply the laser bias current, ramping the current at 10 mA/sec up to I_bias
  4. For the delay line measurement, scan the laser current until there is a beat note near 80 MHz
    • The 1611 may be saturating. If harmonics of the main beat note are visible,
      • iteratively reduce the laser current and adjust the TEC setpoint while keeping the beat note near 80 MHz
      • Since the laser power depends more strongly on current, this will reduce the beat note power. Do so until no harmonics are visible on the spectrum analyzer.
      • Empirically, I found this to be true when the total power on the photodiode was under 500 uW

Notes about the lasers...

  • If it is required to set the bias current and TEC current simultaneously, first tune the laser to several degrees below the set point in the hysteresis free region, then slowly increase the TEC temperature to the setpoint
  • Ensure that the TEC control loop is restricted from entering the mode hop region, since this will put us on the lower hysteresis curve
  • For the North laser (SN: 104978), T_set is 25 C, corresponding to about 9.9-10.1 kOhm. Operated at the recommended I_bias of 150 mA, I observed the next mode hop at 8.971 kOhm, so the TEC control should be kept above that level (operated at lower bias currents, the hop has been observed at lower values of TEC resistance).
  • For the South laser (SN: 104987), T_set is 23 C, corresponding to about 10.8-11.04 kOhm.

 Measuring current noise

  1. Turn off the laser by
    1. Ramp down then turn off LD current
    2. Turn off TEC
    3. Turn off driver box
  2. Remove DB9 at the driver box, and add a DB9 breakout between the driver and the cable going to the laser.
    • The voltage across the diode is between pins 3 and 7
  3. Send the voltage across pins 3 and 7 to a floating-input DC coupled SR560, and the output of the SR560 to an oscilloscope
  4. Turn on the laser by
    1. Turn on driver box
    2. Turn on TEC
    3. Turn on LD current by
      1. With the current off, turn the current setpoint to 0
      2. Turn on the LD current
      3. Ramp the LD current by 10 mA/s up to the desired setpoint
  5. Note the voltage across the diode at 2-3 different values of bias current near the operating point. This is the voltage to current calibration.
  6. Turn off the laser as above
  7. Switch the SR560 to AC coupled mode by
    1. Disconnect the BNC from the SR560 input
    2. switch the preamplifier to AC coupled mode
    3. Change the gain to G=100
    4. reconnect the BNC to SR560 input.
  8. Turn on the laser as above
  9. Record the output of the SR560 with a spectrum analyzer. Note the settings of the SR560, and use that and the calibration above to convert the spectrum from V/rtHz to A/rtHz of the current driver.
    • The delay line frequency discriminator calibration can then be used to convert from A/rtHz to Hz/rtHz
  10. Return the system to its original state by
    1. Turn off laser as above
    2. Remove DB9 breakout and reattach DB9 cable directly to LD driver
Laser TEC setpoint diode setpoint (I_0) Voltage at I_0 diode current (I_1) Voltage at I_1 diode current (I_2) Voltage at I_2 R_sense

Rio N

(ITC510)

9.548
 kOhm
 95.9
 mA
-1.84 V 100.4 mA -1.89 V 92.5
 mA
-1.82 V 8.98 Ohm
Rio S (ITC502) 9.515 kOhm 122.92 mA -2.15 V 119.05 mA -2.11 V 126.53 mA -2.18 V 9.37 Ohm

Notes on current noise measurements:

  • SR560 settings: G=100, AC coupled, 30 kHz LP with 6 db/oct, battery powered
  • Updated on May 14 with S laser drive (ITC502) current noise. Measured at 123.43 mA.
  2736   Tue May 11 13:13:46 2021 aaronLab InfrastructureGeneralproperty tags

I confirmed the inventory of LIGO equipment in cryo lab. Liz requested photos of these tags to clear things up.

 

Attachment 1: IMG_0589.jpeg
IMG_0589.jpeg
Attachment 2: IMG_0586.jpeg
IMG_0586.jpeg
Attachment 3: IMG_0585.jpeg
IMG_0585.jpeg
Attachment 4: IMG_0584.jpeg
IMG_0584.jpeg
Attachment 5: IMG_0583.jpeg
IMG_0583.jpeg
  2735   Tue May 11 13:10:27 2021 aaronLab InfrastructurePurchasesconfirming enclosure dimensions

Shruti and I confirmed that the dimensions for the enclosure make sense. We measured

  • 100" between floor and bottom of lights
  • 68" between the two lights above PSOMA table
  • 105" floor to ceiling beam
  • ~27" tabletop to corner of drop ceiling. Maybe 5-10" to spare here if we wanted to move the table over
  • end of table to beam is 42-43". Based on our independent understandings of the drawings, the beam will not overlap with the HEPA FFU after installing the enclosure.
  • Beam is 11" wide

Note to self: line the corner of the drop ceiling with some foam.

ELOG V3.1.3-